OpenStudy (anonymous):

find the reciprocal of the number 4/7

OpenStudy (anonymous):

what is a recprical?

OpenStudy (anonymous):

A reciprocal is the inverse of the number.

OpenStudy (anonymous):

so it would be 7/4?

OpenStudy (anonymous):

yes

OpenStudy (bibby):

1/4/7

OpenStudy (anonymous):

oh ok so does this apply to any quetion like this?

OpenStudy (anonymous):

question

OpenStudy (anonymous):

Pretty much. When you have a whole number. You do 1 divided by the number. Example: find the reciprocal of 8. It is 1/8

OpenStudy (anonymous):

oh ok so what Bibby saying? i thought it was just 7/4?

OpenStudy (anonymous):

It is 7/4

OpenStudy (bibby):

1/(4/7) = 1*(7/4) = 7/4

OpenStudy (bibby):

yeah

OpenStudy (anonymous):

oh ok i got it now. THank you guys

OpenStudy (anonymous):

Glad to help